A peach breeder has an orchard of peach trees with an average fruit mass of 30 g. The breeder selects the top 10% of plants to reproduce for the next generation. The selected plants have an average fruit mass of 45 g. The next generation has an average fruit mass of 36 g. What is the narrow-sense heritability for fruit weight in this orchard?

____________

The breeder is also wanting to increase sugar and vitamin D content, and she has measured the following variances for these traits:

Variance Sugar Vitamin D
VP 36 80
VG 12 40
VA 8 12

Which trait will respond best to selection imposed by the breeder?

__________
Answer #1

R=h2 X s

R = 36 - 30 = 6
S = 45-30 = 15

h2 = R/S = 6/15 = 0.4


2. h2 = Va / Vp

Sugar h2 = 8/36 = 0.22

Vitamin D h2 = 12/80 = 0.15

Higher h2 is selected. So, sugar